OpenStudy (anonymous):

How many molecules are present in 936 g of glucose (C6H12O6)? (The molar mass of glucose = 180.16 g/mol.)

OpenStudy (chmvijay):

180.16 g of glucose contains 6.023*10^23 molecules of glucose 936 gram of glucose contains = 6.023*10^23 *936 /180.16 = find it

OpenStudy (anonymous):

\[n=\frac{ N }{ L }\] since \[n=\frac{ m }{ M }\] what you now get is, \[\frac{ m }{ M }=\frac{ N }{ L }\]

OpenStudy (anonymous):

n=no. of moles N=Number of particles (atoms, molecules, ions....) L=Avogadro's constant (6.02*10^23) m=mass M=Molar mass

OpenStudy (anonymous):

plug in the values you have and find N (number of molecules).